OpenStudy (anonymous):

Find the exact value of the remaining trigonometric functions of theta: cos(theta)=-3/5, theta in quadrant 3

OpenStudy (paxpolaris):

in quadrant 3:

OpenStudy (paxpolaris):



OpenStudy (paxpolaris):

so what are \(\sin(\theta)\), \(\tan(\theta)\) ....

OpenStudy (paxpolaris):

\[\sin(\theta)={-4\over5}\]\[\tan(\theta)={-4\over-3}=\frac43\] .... and if you need find cot, sec or cosecant too, you can do that

OpenStudy (anonymous):

How did you get -4?